# Mokav: Execution-driven Differential Testing with LLMs

### v1

@@
Abstract.
It is essential to detect functional differences in various software engineering tasks, such as automated program repair, mutation testing, and code refactoring. The problem of detecting functional differences between two programs can be reduced to searching for a difference exposing test (DET): a test input that results in different outputs on the subject programs. In this paper, we propose Mokav, a novel execution-driven tool that leverages LLMs to generate DETs. Mokav takes two versions of a program (P and Q) and an example test input. When successful, Mokav generates a valid DET, a test input that leads to different outputs on P and Q. Mokav iteratively prompts an LLM with a specialized prompt to generate new test inputs. At each iteration, Mokav provides execution-based feedback regarding previously generated tests until the LLM produces a DET. We evaluate Mokav on 1,535 pairs of Python programs collected from the Codeforces competition platform and 32 pairs of programs from the QuixBugs dataset. Our experiments show that Mokav outperforms the state-of-the-art, Pynguin and Differential Prompting (ASE 2023), by a large margin. Mokav can generate DETs for 81.7% (1,255/1,535) of the program pairs in our benchmark (versus 4.9% for Pynguin and 37.3% for Differential Prompting). We demonstrate that all components in our system, including the iterative and execution-driven approaches, contribute to its high effectiveness.
@@ Abstract.
copyright: none
1. Introduction
In differential testing, two versions of a program are given (P & Q), and a test input is searched for such that P & Q produce different outputs (Gulzar et al., 2019). This test exposes a functional difference between P & Q, thus is called a difference exposing test (DET). DETs are useful in various software development tasks, such as for explaining changes (Castellano et al., 2022), detecting anomalies and bugs (Jarman et al., 2020a; Gulzar et al., 2019), analyzing refactorings (Daniel et al., 2007), etc. Generating a DET is significantly more challenging than random test generation because it requires exploring the vast input space of P & Q for the rare inputs that trigger the functional differences (McKeeman, 1998).
Existing techniques to find difference exposing tests limit the search scope using symbolic execution (Rutledge and Orso, 2022), type aware mutation (Liu et al., 2024c), and code coverage optimization with genetic algorithms (Da Silva et al., 2020). These techniques do not leverage recent advanced methods for understanding P & Q’s semantics to guide the search for DETs. Specifically, large language models (LLMs) have shown strong performance in program comprehension (Yuan et al., 2023). This makes LLMs an interesting candidate for difference exposing test search. Previous studies have employed LLMs to conduct various software testing tasks, such as test completion (Nie et al., 2023), test input generation (Xia et al., 2024), test oracle generation (Dinella et al., 2022), unit test generation (Schäfer et al., 2023), and GUI testing (Liu et al., 2024a). To the best of our knowledge, there is no prior work on leveraging LLMs for differential testing.
@@ 1. Introduction
We evaluate Mokav on two datasets: 32 pairs of Python programs in the widely used QuixBugs (Lin et al., 2017) dataset, and our own crafted dataset C4DET. We create C4DET by carefully selecting 1,535 pairs of Python programs from Codeforces [^1] https://codeforces.com/ competitions, with a guaranteed functional difference. Our evaluation results show that Mokav outperforms our two strong baselines, Pynguin (Lukasczyk and Fraser, 2022) and Differential Prompting (DPP(Li et al., 2023). Mokav generates a DET for 81.7% (1,255/1,535) of the pairs in C4DET, while Pynguin and DPP generate a DET for 4.9% and 37.3% of the pairs, respectively. This promising result indicates that Mokav effectively leverages LLMs for differential testing. Mokav is the first LLM-based differential testing, and we provide a publicly available implementation for future research.
@@
  • We introduce a large, carefully crafted dataset C4DET, collected from Codeforces competitions, to study differential testing in Python. C4DET contains 1,535 program pairs with guaranteed functional differences. We expect this dataset to facilitate future work on differential testing, with or without LLMs.
@@
  • We conduct an ablation study of Mokav demonstrating the impact of each prompt component on its effectiveness. We also show that different LLMs (GPT-3.5, GPT-4o, and CodeLlama) can be successfully plugged into Mokav.
2. Problem Statement
In many software engineering tasks, developers face multiple versions of a program, having to reason about their functional difference or equivalence. For example, developers want to assess equivalence after refactoring. Another example is when developers need to ensure the functional difference of a mutant in mutation testing. A difference exposing test (DET) is a test input on which two given programs function differently at runtime (Taneja and Xie, 2008a), it is an existential, concrete proof of functional difference.
@@ 2. Problem Statement
Generating a DETs is hard. The input space of the given programs can be excessively large, which makes it impossible to check the entire input space for finding a DET. Consequently, generating a DET requires an efficient exploration of the input space, based on guesses of potential functional differences between the given programs. We note that the execution of programs under analysis for each sample inputs can also take time (Pizzoleto et al., 2019).
Hence, an efficient DET generation tool must have two qualities: (1) leverage as much static information as possible for comprehending the behavior differences of given programs, and (2) perform a targeted input exploration to narrow down the input space before paying the price of execution. Both problems have been studied in recent years (VieiraB et al., 2022; Schoofs et al., 2022), but remain open.
Note that per Definition 2.1, a DET detects functional differences between P and Q, meaning they produce different outputs for a given input. In this work, we consider functional difference to be distinct from the execution difference. P and Q have an execution difference if, for a given input, there is a variable v used in both P and Q where the set of values v takes on in P is different from the set of values v takes on in Q. In essence, functional difference relates to external differences between P and Q that can be identified by an external observer; in contrast, execution difference pertains to internal differences between P and Q that are detected solely through monitoring the internal states of P and Q. The goal of this paper is to design a novel, effective approach for DET generation.
3. Approach
3.1. Overview
Figure 1. Overview of Mokav. Mokav exposes functional differences between two programs P and Q by feeding LLMs with execution feedback.
Figure 2. The initial prompt (pr0) used in Mokav.
@@ 35 Q: "SEGMENT"
Our novel technique of generating DETs with LLMs employs two powerful features of LLMs. First, Mokav exploits the LLM’s ability to explore the input space for a given program. As previous work demonstrates, given appropriate information, LLMs can generate a high number of diverse and novel test inputs (Schäfer et al., 2023). This enables us to employ LLMs for exploring different parts of the input space of P and Q. Second, we leverage the LLM’s ability to understand the semantic differences between closely related programs. Recent studies show that advanced LLMs can accurately detect inconsistencies between programs and their descriptions (Li and Shin, 2024). This suggests that such LLMs can also accurately understand the differences between P and Q. We take advantage of this strength of LLMs to capture the functional differences between P and Q. Providing execution data to the LLM is a major part of Mokav’s design to improve LLMs’ understanding of P and Q.
3.2. Input
@@
3.3. Mokav Initial Phase
@@
(1) Example test output and execution difference: The runtime analyzer unit of Mokav executes the example test on P and Q to obtain the corresponding output and identify execution differences. Note that the output for the example test must be the same on both versions. If not, the generation is aborted. To identify the execution differences the runtime analyzer takes three steps as follows. First, it finds the variables that are common between P and Q. Next, it uses spotflow [^3] https://github.com/andrehora/spotflow to collect variable values during the execution of P and Q. Finally, the runtime analyzer identifies the first time a common variable gets a value in one of the versions, while that variable value never occurs in the other version. Such variable values are called unique variable values (Etemadi et al., 2023), and the runtime analyzer reports them as the detected execution difference. Note that depending on the example test, the runtime analyzer may be unable to detect a unique variable value for either of the versions. In this case, that part of the data is excluded from the prompt. The output for an example test and the detected execution differences clarify the behavior of the programs, which is crucial for DET generation.
(2) Program descriptions: Mokav prompts the LLM to generate natural language descriptions for P and Q. Recent studies show advanced LLMs are able to understand inconsistencies between programs and their natural language descriptions (Li and Shin, 2024). This suggests that adding the generated descriptions to pr0 can help the LLM to detect behavioral differences between P and Q. The prompt Mokav uses for description generation is a single line message “What is the intention of this code?”, followed by the program.
The pr0 follows the structure shown in Figure 2. The main part of the prompt consists of (1) the P and Q versions of the program, (2) the description of P and Q, (3) the example test and its corresponding output, (4) the execution difference detected while running the example test, and (5) text to specify the kind of response that is expected. For the latter, the expected format of the generated test is presented. Then, the prompt defines what is a difference exposing test. To ensure the LLM understands the definition, it is written in formal terms as well as natural language. The formal definition (“P(inputdate)!=Q(inputdata”) helps the model relate the DET generation task with the given P and Q programs.
@@
3.4. Mokav Iteration Phase
@@
pr1 has three parts. The first part is the detected execution difference. If the runtime analyzer does not detect an execution difference, this part is excluded for pr1. The second part informs the LLM that P and Q produce the same output for the previously generated test, while we expect a test with different outputs on P and Q. The last part of the pr1 directly asks for another test.
At the end of the iteration phase, the result is either a generated test that exposes the functional difference or a failure message that the maximum number of iterations is reached.
3.5. Outcome of Mokav
The generated difference exposing test is presented to the user in the form of a unit test class. This test case imports P and Q as two functions fp and fq. Then a test case method is passes the generated DET to fp and fq and asserts that their outputs are not equal (for example, see the outcome section of Figure 1).
3.6. Implementation
@@
4. Experimental Methodology
4.1. Research Questions
@@
  • RQ1 (effectiveness): How effective is Mokav for generating difference exposing tests? We run Mokav on QuixBugs (Lin et al., 2017) and a large dataset of pairs of programs with functional differences collected from Code4Bench (Majd et al., 2019). We assess the effectiveness of Mokav in DET generation and compare it against two baselines: (1) a state-of-the-art search-based test generation tool, called Pynguin (Lukasczyk and Fraser, 2022), and (2) DPP (Li et al., 2023).
@@
4.2. Datasets
We employ two datasets in our experiments. First, QuixBugs (Lin et al., 2017) is a dataset of 40 pairs of programs in Python that implement common programming problems, such as the greatest common divisor. For each of the 40 pairs of programs in QuixBugs, one version is buggy and the other fixed. There is at least one input on which the buggy version either faces runtime error or has a return value different from the fixed version. This guarantees the functional difference between the two versions, with a significant functional overlap. This means that a DET is not easy to find. In this paper, we consider 32 program pairs from QuixBugs. For these pairs, QuixBugs contains a test that produces the same output on the buggy and fixed versions. This test should be used as the example test in Mokav’s workflow.
As a second dataset used in our evaluation, we create a curated dataset of pairs of programs, which we call C4DET. C4DET contains pairs of similar programs with guaranteed functional differences. We create C4DET by carefully collecting and curating submissions to the Codeforces code competition website. We use Code4Bench (Majd et al., 2019) to extract the required information for C4DET. Code4Bench contains problems, authors, submissions, and tests in Codeforces. Each submission is labeled with a verdict, which shows whether it passes the tests for the corresponding problem, gives a wrong answer, causes runtime error, causes timeout, or causes memory limit. C4DET consists of a subset of the programs and their corresponding problems and test cases in Code4Bench. These programs are selected as follows: for every author $a$ and every problem $pr$ in Codeforces, we add $\langle P,Q\rangle$ to C4DET iff they meet the following conditions:
  1. (1)
    $P$ and $Q$ are both Python submissions by $a$ for $pr$.
  2. (2)
@@ 4.2. Datasets
  6. (6)
    There are at least two submissions in C4DET for $pr$ that pass all the tests. We add this condition to have at least two reference versions to be used for oracle assessment (Liu et al., 2024b).
The programs in Code4Bench are Python scripts that read inputs from stdin and print outputs to stdout. As many traditional test generation tools, such as Pynguin (Lukasczyk and Fraser, 2022), generate tests for functions, we transform all programs to their function-based versions in C4DET. LABEL:lst:function_based shows an example of this transformation. The source in Code4Bench is a Python script that reads a string (line 3) and prints the output (lines 6 and 8). The transformed version is a Python function that receives a list of arguments in *args (line 11) and creates a return_list (line 12) that is returned as the output of the function (line 21). The first input is read from args (line 15) and the outputs are added to the return_list at lines 18 and 20. This transformed version keeps the order of inputs and outputs as well as the functionality of the original source. At the same time, the transformed version is a function-based program suited for test generation without input/output.
Overall, C4DET contains 1,535 pairs of programs for 221. The longest pair of programs in C4DET has 4,242 tokens when tokenized by the gpt-3.5 tokenizer. Many advanced LLMs, such as gpt-3.5, have a 16,000 token limit. Thus, the maximum number of tokens in C4DET program pairs is low enough to be below this limit and leave space for other components of our prompt (see Figure 2). The median number of tokens in C4DET pairs of programs is 248. This is a reasonable number of tokens for a Python function. It is also close to the median number of tokens in Code4Bench, meaning that C4DET is representative of programs in online competitions, which are used in many software engineering research projects (Majd et al., 2019; Hendrycks et al., 2021).

### v2

@@
Abstract
It is essential to detect functional differences between programs in various software engineering tasks, such as automated program repair, mutation testing, and code refactoring. The problem of detecting functional differences between two programs can be reduced to searching for a difference exposing test (DET): a test input that results in different outputs on the subject programs. In this paper, we propose Mokav, a novel execution-driven tool that leverages LLMs to generate DETs. Mokav takes two versions of a program (P and Q) and an example test input. When successful, Mokav generates a valid DET, a test input that leads to provably different outputs on P and Q. Mokav iteratively prompts an LLM with a specialized prompt to generate new test inputs. At each iteration, Mokav provides execution-based feedback from previously generated tests until the LLM produces a DET. We evaluate Mokav on 1,535 pairs of Python programs collected from the Codeforces competition platform and 32 pairs of programs from the QuixBugs dataset. Our experiments show that Mokav outperforms the state-of-the-art, Pynguin and Differential Prompting, by a large margin. Mokav can generate DETs for 81.7% (1,255/1,535) of the program pairs in our benchmark (versus 4.9% for Pynguin and 37.3% for Differential Prompting). We demonstrate that the iterative and execution-driven feedback components of the system contribute to its high effectiveness.
keywords:
@@ Abstract
journal: Journal of Systems and Software
1 Introduction
In differential testing, two versions of a program are given (P & Q), and a test input is searched for such that P & Q produce different outputs [1]. This test exposes a functional difference between P & Q, thus is called a difference exposing test (DET). DETs are useful in various software development tasks, such as for explaining changes [2], detecting anomalies and bugs [3, 1], analyzing refactorings [4]. Generating a DET is significantly more challenging than random test generation because it requires exploring the vast input space of P & Q for the rare inputs that trigger the functional differences [5].
Existing techniques to find difference exposing tests limit the search scope using symbolic execution [6], type aware mutation [7], and code coverage optimization with genetic algorithms [8]. These techniques do not leverage recent advanced methods for understanding P & Q’s semantics to guide the search for DETs. Specifically, large language models (LLMs) have shown strong performance in program comprehension [9]. This makes LLMs an interesting candidate for difference exposing test search. Previous studies have employed LLMs to conduct various software testing tasks, such as test completion [10], test input generation [11], test oracle generation [12], unit test generation [13], and GUI testing [14]. To the best of our knowledge, there is no prior work on leveraging LLMs for differential testing.
@@ 1 Introduction
We evaluate Mokav on two datasets: 32 pairs of Python programs in the widely used QuixBugs [15] dataset, and our own crafted dataset C4DET. We create C4DET by carefully selecting 1,535 pairs of Python programs from Codeforces [^3] https://codeforces.com/ competitions, with a guaranteed functional difference. Our evaluation results show that Mokav outperforms our two strong baselines, Pynguin [16] and Differential Prompting (DPP[17]. Mokav generates a DET for 81.7% (1,255/1,535) of the pairs in C4DET, while Pynguin and DPP generate a DET for 4.9% and 37.3% of the pairs, respectively. This promising result indicates that Mokav effectively leverages LLMs for differential testing. Mokav is the first LLM-based differential testing, and we provide a publicly available implementation for future research.
@@
  • We introduce a large, carefully crafted dataset C4DET, collected from Codeforces competitions, to study differential testing in Python. C4DET contains 1,535 program pairs with guaranteed functional differences. We expect this dataset to facilitate future research on differential testing, with or without LLMs.
@@
  • We conduct an ablation study of Mokav demonstrating the impact of each component on its effectiveness. We also show that different LLMs (GPT-3.5, GPT-4o, and CodeLlama) can be successfully plugged into Mokav.
2 Problem Statement
In many software engineering tasks, developers face multiple versions of a program, having to reason about their functional difference or equivalence. For example, developers want to assess equivalence after refactoring. Another example is when developers need to ensure the functional difference of a mutant in mutation testing. A difference exposing test (DET) is a test input on which two given programs function differently at runtime [18], it is an existential, concrete proof of functional difference.
@@ 2 Problem Statement
Generating a DETs is hard. The input space of the given programs can be excessively large, which makes it impossible to check the entire input space for finding a DET. Consequently, generating a DET requires an efficient exploration of the input space, based on guesses of potential functional differences between the given programs. We note that the execution of programs under analysis for each sample inputs can also take time [19].
Hence, an efficient DET generation tool must have two qualities: (1) leverage as much static information as possible for comprehending the behavior differences of given programs, and (2) perform a targeted input exploration to narrow down the input space before paying the price of execution. Both problems have been studied in recent years [20, 21], but remain open.
Note that per Definition 1, a DET detects functional differences between P and Q, meaning they produce different outputs for a given input. In this work, we consider functional difference to be distinct from the execution difference. P and Q have an execution difference if, for a given input, there is a variable v used in both P and Q where the set of values v takes on in P is different from the set of values v takes on in Q. In essence, functional difference relates to external differences between P and Q that can be identified by an external observer; in contrast, execution difference pertains to internal differences between P and Q that are detected solely through monitoring the internal states of P and Q. The goal of this paper is to design a novel, effective approach for DET generation.
3 Approach
3.1 Overview
Figure 1: Overview of Mokav. Mokav exposes functional differences between two programs P and Q by feeding LLMs with execution feedback.
@@ 35 Q: "SEGMENT"
Our novel technique of generating DETs with LLMs employs two powerful features of LLMs. First, Mokav exploits the LLM’s ability to explore the input space for a given program. As previous work demonstrates, given appropriate information, LLMs can generate a high number of diverse and novel test inputs [13]. This enables us to employ LLMs for exploring different parts of the input space of P and Q. Second, we leverage the LLM’s ability to understand the semantic differences between closely related programs. Recent studies show that advanced LLMs can accurately detect inconsistencies between programs and their descriptions [22]. This suggests that such LLMs can also accurately understand the differences between P and Q. We take advantage of this strength of LLMs to capture the functional differences between P and Q. Providing execution data to the LLM is a major part of Mokav’s design to improve LLMs’ understanding of P and Q.
3.2 Input
@@
3.3 Mokav Initial Phase
Figure 2: The initial prompt (pr0) used in Mokav.
@@ 3.3 Mokav Initial Phase
(1) Example test output and execution difference: The runtime analyzer unit of Mokav executes the example test on P and Q to obtain the corresponding output and identify execution differences. Note that the output for the example test must be the same on both versions. If not, the generation is aborted. To identify the execution differences the runtime analyzer takes three steps as follows. First, it finds the variables that are common between P and Q. Next, it uses spotflow [^5] https://github.com/andrehora/spotflow to collect variable values during the execution of P and Q. Finally, the runtime analyzer identifies the first time a common variable gets a value in one of the versions, while that variable value never occurs in the other version. Such variable values are called unique variable values [23], and the runtime analyzer reports them as the detected execution difference. Note that depending on the example test, the runtime analyzer may be unable to detect a unique variable value for either of the versions. In this case, that part of the data is excluded from the prompt. The output for an example test and the detected execution differences clarify the behavior of the programs, which is crucial for DET generation.
(2) Program descriptions: Mokav prompts the LLM to generate natural language descriptions for P and Q. Recent studies show advanced LLMs are able to understand inconsistencies between programs and their natural language descriptions [22]. This suggests that adding the generated descriptions to pr0 can help the LLM to detect behavioral differences between P and Q. The prompt Mokav uses for description generation is a single line message “What is the intention of this code?”, followed by the program.
The pr0 follows the structure shown in Figure 2. The main part of the prompt consists of (1) the P and Q versions of the program, (2) the description of P and Q, (3) the example test and its corresponding output, (4) the execution difference detected while running the example test, and (5) text to specify the kind of response that is expected. For the latter, the expected format of the generated test is presented. Then, the prompt defines what a difference exposing test is. To ensure the LLM understands the definition, it is written in formal terms as well as natural language. The formal definition (“P(inputdate)!=Q(inputdata”) helps the model relate the DET generation task with the given P and Q programs.
@@
3.4 Mokav Iteration Phase
@@
At each iteration, Mokav invokes the LLM as a continuation of the initial prompt and previous iterations’ prompts. This means that the request to LLM in the ith iteration $it_{i}$ also contains the prompts from the $it_{1},it_{2},...,it_{i-1}$ iterations and the initial prompt. Therefore, the LLM leverages information regarding all previous attempts as well as the example test and descriptions at each iteration.
The pr1 prompt has three parts. The first part is the detected execution difference. If the runtime analyzer does not detect an execution difference, this part is excluded for pr1. The second part informs the LLM that P and Q produce the same output for the previously generated test, while we expect a test with different outputs on P and Q. The last part of the pr1 directly asks for another test.
At the end of the iteration phase, the result is either a generated test that exposes the functional difference or a failure message that the maximum number of iterations has been reached.
3.5 Outcome of Mokav
The generated difference exposing test is presented to the user in the form of a unit test class. This test case imports P and Q as two functions fp and fq. Then a test case method passes the generated DET to fp and fq and asserts that their outputs are not equal (for example, see the outcome section of Figure 1).
3.6 Implementation
@@
4 Experimental Methodology
4.1 Research Questions
@@
  • RQ1 (effectiveness): How effective is Mokav for generating difference exposing tests? We run Mokav on QuixBugs [15] and a large dataset of pairs of programs with functional differences collected from Code4Bench [24]. We assess the effectiveness of Mokav in DET generation and compare it against two baselines: (1) a state-of-the-art search-based test generation tool, called Pynguin [16], and (2) DPP [17].
@@
4.2 Datasets
We employ two datasets in our experiments. First, QuixBugs [15] is a dataset of 40 pairs of programs in Python that implement common programming problems, such as the greatest common divisor. For each of the 40 pairs of programs in QuixBugs, one version is buggy and the other is fixed. There is at least one input on which the buggy version either faces runtime error or has a return value different from the fixed version. This guarantees the functional difference between the two versions, with a significant functional overlap. This means that a DET is not easy to find. In this paper, we consider 32 program pairs from QuixBugs. For these pairs, QuixBugs contains a test that produces the same output on the buggy and fixed versions. This test should be used as the example test in Mokav’s workflow.
As a second dataset used in our evaluation, we create a curated dataset of pairs of programs, which we call C4DET. C4DET contains pairs of similar programs with guaranteed functional differences. We create C4DET by carefully collecting and curating submissions to the Codeforces code competition website. We use Code4Bench [24] to extract the required information for C4DET. Code4Bench contains problems, authors, submissions, and tests in Codeforces. Each submission is labeled with a verdict, which shows whether it passes the tests for the corresponding problem, gives a wrong answer, causes runtime error, causes timeout, or causes memory limit. C4DET consists of a subset of the programs and their corresponding problems and test cases in Code4Bench. These programs are selected as follows: for every author $a$ and every problem $pr$ in Codeforces, we add $\langle P,Q\rangle$ to C4DET iff they meet the following conditions:
  1. 1.
    $P$ and $Q$ are both Python submissions by $a$ for $pr$. $P$ and $Q$ are Python programs that do not use third-party libraries and are fully executable without manual environment setup.
  2. 2.
@@ 4.2 Datasets
  6. 6.
    There are at least two submissions in C4DET for $pr$ that pass all the tests. We add this condition to have at least two reference versions to be used for oracle assessment [25].
The programs in Code4Bench are Python scripts that read inputs from stdin and print outputs to stdout. As many traditional test generation tools, such as Pynguin [16], generate tests for functions, we transform all programs to their function-based versions in C4DET. The transformation modifies input/outputs in two steps. First, it replaces stdin inputs with elements of an arguments array given to the function as *args. Secondly, it adds all the printed strings to newly defined return_list array. The original print statements are removed and the function returns the return_list array. LABEL:lst:function_based shows an example of this transformation. The source in Code4Bench is a Python script that reads a string (line 3) and prints the output (lines 6 and 8). The transformed version is a Python function that receives a list of arguments in *args (line 11) and creates a return_list (line 12) that is returned as the output of the function (line 21). The first input is read from args (line 15) and the outputs are added to the return_list at lines 18 and 20. This transformed version keeps the order of inputs and outputs as well as the functionality of the original source. At the same time, the transformed version is a function-based program suited for test generation without input/output.
Overall, C4DET contains 1,535 pairs of programs for 221 problems. The longest pair of programs in C4DET has 4,242 tokens when tokenized by the gpt-3.5 tokenizer. Many advanced LLMs, such as gpt-3.5, have a 16,000 token limit. Thus, the maximum number of tokens in C4DET program pairs is low enough to be below this limit and leave space for other components of our prompt (see Figure 2). The median number of tokens in C4DET pairs of programs is 248. This is a reasonable number of tokens for a Python function. It is also close to the median number of tokens in Code4Bench, meaning that C4DET is representative of programs in online competitions, which are used in many software engineering research projects [24, 26].